5.a boy has a certain number of sweets. if he eats 16 a day they will last him 2 days longer than if he eats 18 a day. how many sweets has he ?
Found 2 solutions by mananth, fatima2:
Answer by mananth(16949)

a boy has a certain number of sweets. if he eats 16 a day they will last him 2 days longer than if he eats 18 a day. how many sweets has he ?
...
let him have x sweets
..
16 /day
number of days it lasts =x/16
..
if he eats 18/day
number of days = x/18
...
x/16 -x/18 = 2
...
LCD= 18*16
(18x-16x)/18*60=2
multiply by 18*60
18x-16x=2*18*16
2x=2*18*16
/2
x=16*18 =288 sweets
..
m.ananth@hotmail.ca

Answer by fatima2(1)

d=days
take him two days longer=d-2
16days
18(days -2)
16d= 18(d-2)
16d= 18d-36
18d-16d=36
2d=36
d=36/2
d=18
days= 18
=18(d-2)
=18(18-2)
=18(16)
=288